stuck. after (3u-1)^2... becomes u=1/3... to finish i must solve x+3/x+2=1/3
Answer by Eva18(13)

x+3/x+2=1/3.
cross multiply
3(x+3)=1(x+2)
3x+9=x+2
like terms together
3x-x=2-9
2x=-7
divide both sides by 2
x=-3.5.